Find the no of values of n for which 2^8 + 2^11 + 2^n is a perfect square

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
 
2^8 + 2^11 + 2^n 
= 2^8 ( 1 + 2^3 + 2^(n-8) )
= 2^8 (9 + 2^(n-8) )
 
Now for this to be a perfect square,
 
9 + 2^(n-8) sbould be a perfect square.
 
Hence 9 + 16 = 25 is a perfect square.
 
2^(n-8) = 16 = 2^4
n-8 = 4
n = 12